# Acoustic Non-Stationarity Objective Assessment with Hard Label Criteria for Supervised Learning Models

### v1

@@ Abstract
Objective non-stationarity measures are resource intensive and impose critical limitations for real-time processing solutions. In this paper, a novel Hard Label Criteria (HLC) algorithm is proposed to generate a global non-stationarity label for acoustic signals, enabling supervised learning strategies to be trained as stationarity estimators. The HLC is first evaluated on state-of-the-art general-purpose acoustic models, demonstrating that these models encode stationarity information. Furthermore, the first-of-its-kind HLC-based Network for Acoustic Non-Stationarity Assessment (NANSA) is proposed. NANSA models outperform competing approaches, achieving up to 99% classification accuracy, while solving the computational infeasibility of traditional objective measures.
@@ 1 Introduction
Acoustic signals are commonly considered non-stationary across various research domains, including automatic speech recognition (ASR) [1], computational auditory scene analysis (CASA) [2], and speech enhancement (SE) [3, 4]. However, despite the usual assumption, experiments are rarely accompanied of objective assessments, which are essential to validate the hypothesis and evaluate strategies under different degrees of temporal and spectral variations.
@@ 2.1 Review of the INS Framework
where $\langle\ S(t_{h},\boldsymbol{\cdot})\ \rangle_{z}$ is the spectrogram of section $z$ for local observation window $T_{h}<T/2$ and scale $T_{h}/T\in(0,0.5]$.
@@
Figure 1 depicts the spectrograms, INS values (in green), and stationarity thresholds $\gamma$ (in blue) for three 1.5-second samples from AudioSet [14]. In the first example, the signal is classified as non-stationary for scales $T_{h}/T<0.4$ and stationary otherwise. That is, only segments with duration $T_{h}\geq 0.4T$ are sufficiently similar to the global spectrogram. In the second case, a clear spectral pattern is observed, and the signal is non-stationary for $T_{h}/T<0.2$, indicating that only shorter segments (less than 0.3 seconds) exhibit spectral distributions sufficiently distinct from the global pattern. Finally, in the last example, the spectral energy distribution remains consistent over time, and the signal is stationary across all observable scales.
@@ 2.2 The Hard Label Criteria (HLC)
In an intuitive analysis, the first and last signals of Figure 1 could be globally categorized due to a common INS behavior for most scales. However, that is not the case for the second example, which illustrates the necessity of a global objective assessment criterion for acoustic non-stationarity.
@@ 2.3 NANSA Architecture and Training Criterion
As an additional contribution of this paper, a specialized Network for Acoustic Non-Stationary Assessment (NANSA) is proposed and illustrated in Figure 2. The architecture consists of three modules - an Acoustic Non-Stationary (ANS) Encoder, a Pattern Extractor, and a Pattern Discriminator.
The Short-Time Fourier Transform (STFT) is applied using 512 samples over $20$ ms frames with 50% overlap, at a 16 kHz sampling rate. The resulting spectrogram $S\in\mathbb{R}^{T_{\text{ANS}}\times 257}$ is passed through a fully connected (FC) layer that scales the frequency dimension by a factor $\beta_{FC}$, followed by a ReLU activation and a second FC layer with inverse scaling $1/\beta_{FC}$, which leads to the embedding $E_{ANS}\in\mathbb{R}^{T_{\text{ANS}}\times 257}$. A classification embedding $E_{CLS}$ is appended to $E_{ANS}$. As the INS computation operates on spectrogram segments, unitary temporal patches and positional embeddings are incorporated, as in [19]. Non-stationary patterns are extracted using self-attention. Finally, the probability $P_{ANS}$ is computed from the first output embedding of the previous module. The model is trained using binary cross-entropy loss $\mathcal{L}_{BCE}$, with ground truth labels provided by the HLC labeling function $f_{HLC}$ in Eq. 5.
The full NANSA model employs 11 self-attention layers, each with 3 heads and a 192-dimensional input. Its lightweight variant, NANSALW, uses 4 self-attention layers with 3 heads and a 64-dimensional input, designed specifically for resource-constrained devices.
@@ 3.1 Datasets and Baseline Models
The acoustic non-stationarity classification is designed for supervised learning strategies based on HLC labels. Experiments are conducted using signals from AudioSet [14], DCASE [15] and FSD50K [16]. These datasets are originally designed for acoustic sources, scenes and events classification, composing a diverse collection of audio signals. Standard dataset splits are used for training and evaluation.
@@ 3.2 Implementation Details
Audio signals are segmented into 1.5-second clips, consistent with typical durations in speech and on-device audio applications [21, 22]. The HLC algorithm is configured with $\alpha_{HLC}=10$ and $K=3$ partitions over three temporal regions: $\mathcal{P}_{1}=\{0.006,0.012,0.025\}$, $\mathcal{P}_{2}=\{0.05,0.1,0.2\}$, and $\mathcal{P}_{3}=\{0.3,0.4,0.5\}$. These correspond to short- (9.0–37.5 ms), mid- (75–300 ms), and long-term (400–750 ms) temporal dynamics. Using this setup, HLC labeled 63.9% of all acoustic signals as non-stationary.
For NANSA, the $\beta_{FC}$ multiplier is set to 4, as higher values yielded no significant performance gains and unnecessarily increased model size. All models are trained for 20 epochs using a learning rate of $10^{-4}$ and the Adam optimizer [23].
@@ 4 Results and Discussion
Table 2 presents the classification accuracy, EER, and F1-score for HLC-based acoustic non-stationarity assessment. All baseline models (PANNs, AST, and PaSST) achieve over 90% accuracy, indicating that general-purpose acoustic models are capable of encoding non-stationarity information. Among them, the attention-based AST and PaSST outperform PANNs in both accuracy and F1-score, while achieving lower EER values. However, these improvements come with significantly higher memory and compute (MMAC) costs, which are up to an order of magnitude greater than PANNs.
@@
Table 2 also highlights the impact of removing the ANS Encoder from the proposed models. In this ablation, the raw STFT spectrogram replaces the $E_{ANS}$ embedding. This leads to average EER increases of 10.5% and 12.5% for NANSA and NANSALW, respectively, with the latter being more affected due to its smaller capacity.
@@
Figure 3 shows the ROC curves and corresponding AUCs for each dataset. In all scenarios, NANSA and NANSALW curves are closest to the ideal operating point $(0,1)$. Accordingly, the proposed models achieve the highest AUCs, further confirming their efficacy in non-stationarity classification.
It is remarkable that all acoustic models attained consistent classification results via HLC algorithm and therefore can be adopted as a solution to overcome INS resource intensive issues. In Table 3, it is shown the comparison between the inference time of HLC-trained models with the original INS statistical framework. All models significantly reduce processing time compared to INS, thanks to HLC-based training. Notably, NANSA and NANSALW are approximately 500 and 4000 times faster than the traditional INS approach. The time efficiency gain confirms the effectiveness of non-stationarity assessment for HLC-trained models in both large-scale and resource-constrained devices.

### v2

@@ Abstract
Objective non-stationarity measures are resource intensive and impose critical limitations for real-time processing solutions. In this paper, a novel Hard Label Criteria (HLC) algorithm is proposed to generate global non-stationarity labels for acoustic signals, enabling supervised learning strategies to be trained as stationarity estimators. The HLC is first evaluated on state-of-the-art general-purpose acoustic models, demonstrating that these models capture stationarity information. Furthermore, the first-of-its-kind HLC-based Network for Acoustic Non-Stationarity Assessment (NANSA) is proposed. NANSA models outperform competing approaches, achieving up to 99% classification accuracy, while solving the computational infeasibility of traditional objective measures.
@@ 1 Introduction
Acoustic signals are commonly considered non-stationary across various research domains, including automatic speech recognition (ASR) [1], computational auditory scene analysis (CASA) [2], and speech enhancement (SE) [3, 4]. However, despite the usual assumption, experiments are rarely accompanied by objective assessments, which are essential to validate the hypothesis and evaluate strategies under different degrees of temporal and spectral variations.
@@ 2.1 Review of the INS Framework
where $\langle\ S(t_{h},\boldsymbol{\cdot})\ \rangle_{z}$ is the spectrogram of section $z$ for local observation window $T_{h}\leq T/2$ and scale $T_{h}/T\in(0,0.5]$.
@@
Fig. 1 depicts the spectrograms, INS values (in green), and stationarity thresholds $\gamma$ (in blue) for three 1.5-second samples from AudioSet [14]. In the first example, the signal is classified as non-stationary for scales $T_{h}/T<0.4$ and stationary otherwise. That is, only segments with duration $T_{h}\geq 0.4T$ are sufficiently similar to the global spectrogram. In the second case, a clear spectral pattern is observed, and the signal is non-stationary for $T_{h}/T<0.2$, indicating that only shorter segments (less than 0.3 seconds) exhibit spectral distributions sufficiently distinct from the global pattern. In the final example, the spectral energy distribution remains consistent over time, and the signal is stationary across all observable scales.
@@ 2.2 The Hard Label Criteria (HLC)
In an intuitive analysis, the first and last signals of Fig. 1 could be globally categorized due to a common INS behavior for most scales. However, that is not the case for the second example, which illustrates the necessity of a global objective assessment criterion for acoustic non-stationarity.
@@ 2.3 NANSA Architecture and Training Criterion
As an additional contribution, the specialized Network for Acoustic Non-Stationary Assessment (NANSA) is proposed, which consists of three modules as illustrated in Fig. 2.
In the ANS Encoder, the Short-Time Fourier Transform (STFT) is applied every $20$ ms with 50% overlap, at a 16 kHz sampling rate. Resulting spectrogram $S\in\mathbb{R}^{T_{\text{ANS}}\times 257}$ is processed by two fully connected (FC) layers with scaling factors $\beta_{FC}$ and $1/\beta_{FC}$, separated by a ReLU activation, producing the embedding $E_{ANS}$. A classification embedding $E_{CLS}$ is appended to $E_{ANS}$. The transformer-based Pattern Extractor uses self-attention to model both local and long-range temporal dependencies, enabling robust extraction of non-stationary patterns. Since the INS computation operates on spectrogram segments, unitary temporal patches and positional embeddings are employed [19]. The probability $P_{ANS}$ is obtained from the first output embedding of this module. Training is carried with binary cross-entropy loss $\mathcal{L}_{BCE}$, with ground truth labels provided by the $f_{HLC}$ function in (5).
The full NANSA model employs 11 self-attention layers, each with 3 heads and a 192-dimensional input. Its lightweight variant, NANSALW, uses 4 self-attention layers with 3 heads and a 64-dimensional input, targeting resource-constrained devices.
@@ 3.1 Datasets and Baseline Models
The acoustic non-stationarity classification is designed for supervised learning strategies based on HLC labels. Experiments are conducted using signals from AudioSet [14], DCASE [15] and FSD50K [16]. These datasets are originally intended for acoustic sources, scenes and events classification, comprising a diverse collection of audio signals. Standard dataset splits are used for training and evaluation.
@@ 3.2 Implementation Details
Fig. 2: The NANSA model diagram.
Audio signals are segmented into 1.5-second clips, consistent with typical durations in speech and on-device audio applications [21, 22]. Experiments are carried with HLC algorithm configured for $K=3$ regions $\mathcal{T}_{k}$ to capture short-, mid-, and long-term temporal dynamics. The first two regions use approximately geometric progressions, with $\mathcal{T}_{1}=\{0.006,0.012,0.025\}$ (9.037.59.0–37.5 ms) and $\mathcal{T}_{2}=\{0.05,0.1,0.2\}$ (7530075–300 ms), while $\mathcal{T}_{3}=\{0.3,0.4,0.5\}$ (400750400–750 ms) is linear. This design is a reflection of higher observation scales exhibiting slower non-stationarity variations, as illustrated in Fig. 1. Additionally, $\alpha_{HLC}$ is conservatively set to $10$, i.e., HLC defines non-stationarity within a region by imposing a one-order-of-magnitude higher threshold across most INS observable scales. All models are trained for 20 epochs, learning rate of $10^{-4}$ and Adam optimizer [23].
@@ 4 Results and Discussion
Table 2 presents the classification accuracy, EER, and F1-score for HLC-based acoustic non-stationarity assessment. All baseline models (PANNs, AST, and PaSST) achieve over 90% accuracy, indicating that general-purpose acoustic models are capable of capturing non-stationarity information. Among them, the attention-based AST and PaSST outperform PANNs in both accuracy and F1-score, while achieving lower EER values. However, these improvements come with significantly higher memory and compute (MMAC) costs, which are up to an order of magnitude greater than PANNs.
@@
For the ablation study, the impact of the ANS Encoder module is also summarized in Table 2. Preliminary experiments explored $\beta_{FC}\in\{1.5,2,4,6\}$, with $\beta_{FC}=4$ selected due to slightly better performance. Removing the ANS Encoder increases average EER by 10.5% and 12.5% for NANSA and NANSALW, respectively, with the latter being more affected due to its smaller model capacity.
@@
Fig. 3 shows the ROC curves and corresponding AUCs for each dataset. In all scenarios, NANSA and NANSALW curves are closest to the ideal operating point $(0,1)$. Accordingly, the proposed models achieve the highest AUCs, further confirming their efficacy in non-stationarity classification.
It is remarkable that all acoustic models attained consistent classification results via HLC algorithm and therefore can be adopted as a solution to overcome INS resource intensive issues. In Table 3, it is shown the comparison between the inference time of HLC-trained models with the original INS statistical framework. All models significantly reduce processing time compared to INS, thanks to HLC-based training. Notably, NANSA and NANSALW are approximately 500 and 4000 times faster than the traditional INS approach. The time efficiency gain confirms the effectiveness of non-stationarity assessment for HLC-trained models in both large-scale and resource-constrained devices. One drawback of HLC is the loss of scale-relative INS information. As future research, we intent to overcome this with multi-task learning strategies.